sqrt(4 - x) = (4 - x)
= (4 + x) answer
Is this correct?
Thanks

Answer by checkley77(12844)

The square root of a number (4-x) can only equal the number(x-4) IF the number =1.
So x=3 in this case.
Proof:
sqrt(4 - x) = (4 - x)
sqrt(4-3)=(4-3
sqrt1=1
------------------------------
(4+x) is not the answer.